# The Intermediate Value Theorem for Linear Transformations

### v1

@@ Abstract.
If a real-valued function is continuous on a real interval and it takes on two different values, then it will also take any value in between those two, by the Intermediate Value Theorem. It is not immediately clear what would be a natural generalization for functions whose domain and range are in higher-dimensional Euclidean spaces. In this article, we analyze this problem, by first arriving at what we think is the appropriate question to ask, and then restricting to linear transformations. Then we show a class of linear transformations for which we can give a natural generalization of the Intermediate Value Theorem. It turns out that this characterization has been studied before, both in numerical analysis and linear programming.
2020 Mathematics Subject Classification:
15A24, 15A39, 15B48
The author wishes to thank Levent Ulku, who approached the author with a question that led the author to think about the topic of this paper. Also, the author thanks Blair Madore and Edgar Possani for their comments on a first version of this manuscript, which significantly improved the exposition. This work is partially supported by the Asociación Mexicana de Cultura A.C
1. Introduction
@@
Now, what happens if the domain is a subset of $\mathbb{R}^{n}$? Suppose that we have a continuous function $f:D\subseteq\mathbb{R}^{n}\to\mathbb{R}$ and assume $D$ is path-connected (i.e., any two points in $D$ can be joined by a curve contained in $D$). Clearly, if the function $f$ takes the values $c,d\in\mathbb{R}$ , then it must take any value between $c$ and $d$. Indeed, assume $x_{0},x_{1}\in D$ satisfy $f(x_{0})=c$ and $f(x_{1})=d$ and, without loss of generality, that $c<d$. Let $\gamma:[0,1]\to D$ be a continuous function (a “curve”) such that $\gamma(0)=x_{0}$ and $\gamma(1)=x_{1}$. Then $f\circ\gamma$ satisfies the hypothesis of the Intermediate Value Theorem and hence, for every $a\in[c,d]$ there exists $t^{*}\in[0,1]$, such that $(f\circ\gamma)(t^{*})=a$. In other words, there exists $x^{*}:=\gamma(t^{*})\in D$ “between” $x_{0}$ and $x_{1}$ such that $f(x^{*})=a$. Thus, for continuous real-valued functions on path-connected subsets of $\mathbb{R}^{n}$, an Intermediate Value Theorem holds. [^1] In fact, for the more sophisticated reader this is obvious: if $A$ is path-connected, then $A$ is connected. Since $f$ is continuous, then $f(A)$ is a connected set in $\mathbb{R}$ and hence, if it $c$ and $d$ are in the interval $f(A)$, so is any point between $c$ and $d$.
@@
Unfortunately, the answer to this question is also negative. Consider the continuous function $g:\mathbb{R}^{2}\to\mathbb{R}^{2}$ given by $g(x,y)=(x^{2}-y^{2},2xy)$. Observe that $g(0,0)=(0,0)$, that $g(7,1)=(48,14)$, that $(0,0)\leqslant(5,12)\leqslant(48,14)$ and that the only solutions of $g(x,y)=(5,12)$, namely $(3,2)$ and $(-3,-2)$, do not satisfy $(0,0)\leqslant(x,y)\leqslant(7,1)$. Hence, given $(0,0)$ and $(48,14)$, there exists a point, namely $(5,12)$ in between $f(0,0)=(0,0)$ and $f(7,1)=(48,14)$ such that none of the points $(x,y)$ satisfying $f(x,y)=(5,12)$ lie in between $(0,0)$ and $(7,1)$. Hence the answer to the question above is also negative.
@@
2. Linear Transformations: Questions
Let us see what happens if we consider linear functions, so assume $T:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a linear transformation. As it is usual, we will identify such a linear transformation $T$ with its $m\times n$ matrix $A$ (with respect to the canonical bases) so that $T(\mathbf{x})=A\mathbf{x}$; from now on we will speak of the matrix $A$ instead of the linear transformation $T$. We should point out that elements in Euclidean space are thought of as column vectors, so that we can multiply them by matrices; in order to avoid cumbersome spacing in this text if $\mathbf{v}$ is a row vector, we write $\mathbf{v}^{\mathsf{T}}$ to denote its transpose, a column vector.
@@ 2. Linear Transformations: Questions
In order to answer Question 1, we need to decide if sets like $Z$ are contained in the range of $A$. But, with the conditions imposed on $\mathbf{x}$ and $\mathbf{y}$ above, the set $Z$ contains an open set in $\mathbb{R}^{m}$. Since the range of a linear tranformation is a vector subspace, and the only subspace of $\mathbb{R}^{m}$ that contains nontrivial open sets is $\mathbb{R}^{m}$ itself, it follows that if $Z$ is contained in the range of $A$, it must be the case that $A$ is surjective. Hence, unless $A$ is surjective, there will be vectors $\mathbf{z}\in Z$ which are not in the range of $A$. One can get around this issue by requiring some of the coordinates of $\mathbf{x}$ and $\mathbf{y}$ to be equal, but that seems too restrictive. [^3] One can show that for $m\geqslant 2$, if there exists a unique $k^{*}$ such that $x_{k^{*}}=y_{k^{*}}$, then $Z$ is contained in a nontrivial subspace of $\mathbb{R}^{m}$ if and only if $x_{k^{*}}=0$. If $\mathbf{x}$ and $\mathbf{y}$ have $s$ equal coordinates, with $s\geqslant 2$, then $Z$ is contained in an $(m-s)$–dimensional subspace of $\mathbb{R}^{m}$ without any further conditions on the values of the coordinates. Since we do not want to restrict the choices of $\mathbf{x}\leqslant\mathbf{y}$, and since if we require $A$ to be surjective then the answer to Question 1 is trivially affirmative, we will forget about Question 1 and instead focus on Question 2, which we rephrase now for linear transformations.
@@
Let $A$ be an $m\times n$ real matrix. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}\in\mathbb{R}^{n}$ with $\mathbf{x}_{0}\leqslant\mathbf{x}_{1}$; set $\mathbf{y}_{0}:=A\mathbf{x}_{0}$ and $\mathbf{y}_{1}:=A\mathbf{x}_{1}$ and assume that $\mathbf{y}_{0}\leqslant\mathbf{y}_{1}$. Given $\mathbf{y}$ in the range of $A$ with $\mathbf{y}_{0}\leqslant\mathbf{y}\leqslant\mathbf{y}_{1}$, is there a vector $\mathbf{x}\in\mathbb{R}^{n}$ with $A\mathbf{x}=\mathbf{y}$, and $\mathbf{x}_{0}\leqslant\mathbf{x}\leqslant\mathbf{x}_{1}$?
Let us examine the above question. Let
@@
Observe that the matrix $A$ is invertible and hence surjective. Clearly, $A(0,0)^{\mathsf{T}}=(0,0)^{\mathsf{T}}$ and $A(3,1)^{\mathsf{T}}=(15,4)^{\mathsf{T}}$. Now, consider the vector $(10,3)\in\mathbb{R}^{2}$: clearly $(0,0)\leqslant(10,3)\leqslant(15,4)$ but the only solution of the equation $A\mathbf{x}=(10,3)^{\mathsf{T}}$ is $\mathbf{x}=(1,2)$ which does not lie between $(0,0)$ and $(3,1)$. So, Question 3 is false in this case.
@@
Consider the vectors $\mathbf{x}_{0}=(0,0,0)^{\mathsf{T}}$ and $\mathbf{x}_{1}=(1,6,3)^{\mathsf{T}}$. Set $\mathbf{y}_{0}:=A\mathbf{x}_{0}=(0,0)^{\mathsf{T}}$ and $\mathbf{y}_{1}:=A\mathbf{x}_{1}=(4,9)^{\mathsf{T}}$. Consider the vector $\mathbf{y}=(3,1)^{\mathsf{T}}$. Indeed,
@@
00 0 $\leqslant 3-t\leqslant 1,$
00 0 $\leqslant 1-t\leqslant 6,$
00 0 $\leqslant t\leqslant 3.$
and hence there is no vector $(x,y,z)\in\mathbb{R}^{3}$ such that $(0,0,0)\leqslant(x,y,z)\leqslant(1,6,3)$ and $A(0,0,0)^{\mathsf{T}}\leqslant A(x,y,z)^{\mathsf{T}}\leqslant A(1,6,3)^{\mathsf{T}}$.
Hence Question 3 is false for this matrix. But perhaps this example gives a clue to what is a better question to ask. Indeed, notice that if we start with the inequality
@@
So, let us rephrase our question once again, by focusing on the vectors on the range of $A$, instead of starting with vectors in the domain.
@@
Let $A$ be an $m\times n$ real matrix and let $\mathbf{y}_{0}$ and $\mathbf{y}_{1}\in\mathbb{R}^{m}$, both in the range of $A$ such that $\mathbf{y}_{0}\leqslant\mathbf{y}_{1}$. Given $\mathbf{y}$ in the range of $A$ with $\mathbf{y}_{0}\leqslant\mathbf{y}\leqslant\mathbf{y}_{1}$, is it true that there exists $\mathbf{x}_{0}$, $\mathbf{x}$ and $\mathbf{x}_{1}\in\mathbb{R}^{n}$ with $A\mathbf{x}_{0}=\mathbf{y}_{0}$, $A\mathbf{x}=\mathbf{y}$, and $A\mathbf{x}_{1}=\mathbf{y}_{1}$, and such that $\mathbf{x}_{0}\leqslant\mathbf{x}\leqslant\mathbf{x}_{1}$?
As the reader will have observed by now, in the questions above we require $\mathbf{x}_{0}\leqslant\mathbf{x}\leqslant\mathbf{x}_{1}$. The case $\mathbf{x}_{1}\leqslant\mathbf{x}\leqslant\mathbf{x}_{0}$ can be dealt with similarly, as we point out at the end of this article. In the next section, we attempt to answer Question 4.
3. Linear Transformations: Answers
Before trying to answer Question 4, we need to make a couple of observations. The reader will have noticed by now that we have taken $\mathbf{x}_{0}=\mathbf{0}$, the zero vector, in all the examples in the previous section. The reason for that should be clear after the following two propositions. In the first one, we show that when considering Question 4 above for only one inequality, we may assume one side of the inequality equals zero. In the second proposition, we will show that this is in fact equivalent to the inequalities in Question 4.
@@ 3. Linear Transformations: Answers
Let start by proving that (1) implies (2). Assume $\mathbf{0}\leqslant A\mathbf{x}$. Then, $A\mathbf{0}\leqslant A\mathbf{x}$. But (1) implies that there exists vectors $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ such that $A\mathbf{z}_{1}=A\mathbf{0}=\mathbf{0}$, $A\mathbf{z}_{2}=A\mathbf{x}$ and $\mathbf{z}_{1}\leqslant\mathbf{z}_{2}$. Setting $\mathbf{x}^{\prime}:=\mathbf{z}_{2}-\mathbf{z}_{1}\geqslant\mathbf{0}$ we obtain the desired conclusion, since $A\mathbf{x}^{\prime}=A(\mathbf{z}_{2}-\mathbf{z}_{1})=A\mathbf{z}_{2}=A\mathbf{x}$ and $\mathbf{0}\leqslant\mathbf{x}^{\prime}$.
We now show that (2) implies (1). Assume $A\mathbf{x}_{1}\leqslant A\mathbf{x}_{2}$. Then we have that $A(\mathbf{x}_{2}-\mathbf{x}_{1})\geqslant\mathbf{0}$. But then (2) implies that there exists a vector $\mathbf{z}$ such that $A\mathbf{z}=A(\mathbf{x}_{2}-\mathbf{x}_{1})$ and $\mathbf{0}\leqslant\mathbf{z}$. Then, taking $\mathbf{x}_{2}^{\prime}:=\mathbf{x}_{1}+\mathbf{z}$ and $\mathbf{x}_{1}^{\prime}:=\mathbf{x}_{1}$ gives the result, since $\mathbf{x}_{2}^{\prime}-\mathbf{x}_{1}^{\prime}=\mathbf{z}\geqslant\mathbf{0}$, since $A\mathbf{x}_{1}^{\prime}=A\mathbf{x}_{1}$ and since $A\mathbf{x}_{2}^{\prime}=A(\mathbf{x}_{1}+\mathbf{z})=A\mathbf{x}_{2}$, as desired. ∎
It is natural to ask, and perhaps not obvious, if the conditions in the above proposition hold not only for two inequalities, but for three. It turns out that they do hold.
@@
It seems that this definition was first given by Ben-Israel and Greville in [2]. The full characterization of weakly monotone matrices is not easy and it is beyond the scope of this article (much more information can be found in [1, 8]). Nevertheless, we will examine a couple of more examples and give a partial characterization.
Let us examine a previous example again. Let
@@
Let us verify that $A$ is weakly monotone. Indeed, let $\mathbf{x}=(x_{1},x_{2},x_{3})^{\mathsf{T}}$ such that $A\mathbf{x}\geqslant\mathbf{0}$. Then
@@
But clearly we can choose $\mathbf{x}^{\prime}=(x_{1}^{\prime}.x_{2}^{\prime},x_{3}^{\prime})$ with $\mathbf{x}^{\prime}\geqslant\mathbf{0}$ with $x_{1}^{\prime}+x_{3}^{\prime}=x_{1}+x_{3}$ and with $x_{2}^{\prime}+x_{3}^{\prime}=x_{2}+x_{3}$ (for example, take $x_{1}^{\prime}:=x_{1}+x_{3}$, $x_{2}^{\prime}:=x_{2}+x_{3}$ and $x_{3}^{\prime}:=0$) and hence with $A\mathbf{x}=A\mathbf{x}^{\prime}$. Thus $A$ is indeed weakly monotone.
Before we continue talking about weakly monotone matrices, let us state the following definition which was originally given by Collatz [3] for square matrices.
@@
Collatz showed that a square matrix is monotone if and only if it is invertible and its inverse is nonnegative [^4] The anxious reader can try and prove this now: it is not terribly difficult. The patient reader can prove this after seeing the results ahead.
Mangasarian [6] extended the above definition to nonsquare matrices and, furthermore, he showed that a matrix is monotone if and only if $A$ has a nonnegative left inverse. Of course, if $A$ has a left inverse, then $A$ is injective and, in particular, $m\geqslant n$.
In what follows, we intend to give a partial characterization of weakly monotone matrices, using the techniques in [8, 6]. We first start with a famous lemma, whose proof we include here for the sake of completeness. The proof uses the hyperplane separating theorem for convex subsets of Euclidean space (see, for example, [5, p. 510]).
@@
does not contain the vector $\mathbf{b}$. Since $\mathcal{C}$ is a closed convex set there is a hyperplane separating $\mathbf{b}$ from $\mathcal{C}$. That is, there exists $\mathbf{y}\in\mathbb{R}^{m}$, and $\alpha\in\mathbb{R}$ such that the hyperplane $\{\mathbf{w}\in\mathbb{R}^{m}\mid\mathbf{y}^{\mathsf{T}}\mathbf{w}=\alpha\}$ separates $\mathbb{R}^{m}$ in two sides, one containing $\mathbf{b}$ and one containing $\mathcal{C}$; in fact, $\mathbf{y}^{\mathsf{T}}b<\alpha$ and $\mathbf{y}^{T}\mathbf{w}>\alpha$ for all $\mathbf{w}\in\mathcal{C}$.
Observe that since $\mathbf{0}\in\mathcal{C}$, we have $\alpha<0$ and hence $\mathbf{y}^{\mathsf{T}}b<0$. Since for each $\lambda>0$ and for each $k$ the vector $\lambda\mathbf{v}_{k}\in\mathcal{C}$ we have $\mathbf{y}^{\mathsf{T}}\lambda\mathbf{v}_{k}>\alpha$; equivalently, $\mathbf{y}^{\mathsf{T}}\mathbf{v}_{k}>\frac{\alpha}{\lambda}$. Since $\lambda>0$ can be arbitrarily large (and $\alpha<0$), this means that $\mathbf{y}^{\mathsf{T}}\mathbf{v}_{k}\geqslant 0$ for all $k$. But this is equivalent to $\mathbf{y}^{T}M\geqslant\mathbf{0}$. That is, (2) holds.
@@
Let $A$ be an $m\times n$ matrix of full rank. Assume that for all $\mathbf{x}\in\mathbb{R}^{n}$ if we have that $A\mathbf{x}\geqslant\mathbf{0}$ then there exists $\mathbf{x}^{\prime}\in\mathbb{R}^{m}$ such that $\mathbf{x}^{\prime}\geqslant\mathbf{0}$ and $A\mathbf{x}=A\mathbf{x}^{\prime}$. Then,
  • If $m\geqslant n$, then there exists an $n\times m$ matrix $B$ with $BA=I_{n}$ such that $B\geqslant 0$.
  • If $m\leqslant n$, then there exists an $n\times m$ matrix $B$ with $AB=I_{m}$ such that $B\geqslant 0$.

### v2

@@ Abstract.
If a real-valued function is continuous on a real interval and it takes on two different values, then it will also take any value in between those two, by the Intermediate Value Theorem. It is not immediately clear what would be a natural generalization for functions whose domain and range are in higher-dimensional Euclidean spaces. In this article, we analyze this problem, by first arriving at what we think is the appropriate question to ask, and then restricting to linear transformations. It turns out that the matrices that will satisfy an appropriate version of the Intermediate Value Theorem are the so called monotone and weakly monotone matrices, which have applications in numerical approximation of the solutions to systems of linear equations.
1. Introduction.
@@
Now, what happens if the domain is a subset of $\mathbb{R}^{n}$? Suppose that we have a continuous function $f:D\subseteq\mathbb{R}^{n}\to\mathbb{R}$ and assume $D$ is path-connected (i.e., any two points in $D$ can be joined by a curve contained in $D$). Clearly, if the function $f$ takes the values $c,d\in\mathbb{R}$ , then it must take any value between $c$ and $d$. Indeed, assume $x_{0},x_{1}\in D$ satisfy $f(x_{0})=c$ and $f(x_{1})=d$ and, without loss of generality, that $c<d$. Let $\gamma:[0,1]\to D$ be a continuous function (a “curve”) such that $\gamma(0)=x_{0}$ and $\gamma(1)=x_{1}$. Then $f\circ\gamma$ satisfies the hypothesis of the Intermediate Value Theorem and hence, for every $a\in[c,d]$ there exists $t^{*}\in[0,1]$, such that $(f\circ\gamma)(t^{*})=a$. In other words, there exists $x^{*}:=\gamma(t^{*})\in D$ “between” $x_{0}$ and $x_{1}$ such that $f(x^{*})=a$. Thus, for continuous real-valued functions on path-connected subsets of $\mathbb{R}^{n}$, an Intermediate Value Theorem holds. [^1] In fact, for the more sophisticated reader this is obvious: if $A$ is path-connected, then $A$ is connected. Since $f$ is continuous, then $f(A)$ is a connected set in $\mathbb{R}$ and hence, if $c$ and $d$ are in the interval $f(A)$, so is any point between $c$ and $d$.
@@
Unfortunately, the answer to this question is also negative. Consider the continuous function $g:\mathbb{R}^{2}\to\mathbb{R}^{2}$ given by $g(x,y)=(x^{2}-y^{2},2xy)$. Observe that $g(0,0)=(0,0)$, that $g(7,1)=(48,14)$, that $(0,0)\leqslant(5,12)\leqslant(48,14)$, and that the only solutions of $g(x,y)=(5,12)$, namely $(3,2)$ and $(-3,-2)$, do not satisfy $(0,0)\leqslant(x,y)\leqslant(7,1)$. Hence, given $(0,0)$ and $(48,14)$, there exists a point, namely $(5,12)$ in between $f(0,0)=(0,0)$ and $f(7,1)=(48,14)$ such that none of the points $(x,y)$ satisfying $f(x,y)=(5,12)$ lie in between $(0,0)$ and $(7,1)$. Hence the answer to the question above is also negative.
@@
2. Linear Transformations: Questions.
Let us see what happens if we consider linear functions, so assume $T:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a linear transformation. As usual, we will identify such a linear transformation $T$ with its $m\times n$ matrix $A$ (with respect to the canonical bases) so that $T(\mathbf{x})=A\mathbf{x}$; from now on we will speak of the matrix $A$ instead of the linear transformation $T$. Elements in Euclidean space will be thought of as column vectors, so that we can multiply them by matrices. When $\mathbf{v}$ is a row vector, we write $\mathbf{v}^{\mathsf{T}}$ to denote its transpose, a column vector.
@@
In order to answer Question 1, we need to decide if sets like $Z$ are contained in the range of $A$. But, with the conditions imposed on $\mathbf{x}$ and $\mathbf{y}$ above, the set $Z$ contains an open set in $\mathbb{R}^{m}$. Since the range of a linear tranformation is a vector subspace, and the only subspace of $\mathbb{R}^{m}$ that contains nontrivial open sets is $\mathbb{R}^{m}$ itself, it follows that if the set $Z$ is contained in the range of $A$, it must be the case that $A$ is surjective. Hence, unless $A$ is surjective, there will be vectors $\mathbf{z}\in Z$ which are not in the range of $A$. One can get around this issue by requiring some of the coordinates of $\mathbf{x}$ and $\mathbf{y}$ to be equal, but that seems too restrictive. [^3] One can show that for $m\geqslant 2$, if there exists a unique $k^{*}$ such that $x_{k^{*}}=y_{k^{*}}$, then $Z$ is contained in a nontrivial subspace of $\mathbb{R}^{m}$ if and only if $x_{k^{*}}=0$. If $\mathbf{x}$ and $\mathbf{y}$ have $s$ equal coordinates, with $s\geqslant 2$, then $Z$ is contained in an $(m-s)$–dimensional subspace of $\mathbb{R}^{m}$ without any further conditions on the values of the coordinates. Since we do not want to restrict the choices of $\mathbf{x}\leqslant\mathbf{y}$, and since if we require $A$ to be surjective then the answer to Question 1 is trivially affirmative, we will forget about Question 1 and instead focus on Question 2, which we rephrase now for linear transformations.
@@
Let $A$ be an $m\times n$ real matrix. For $\mathbf{x}_{0}$ and $\mathbf{x}_{1}\in\mathbb{R}^{n}$ with $\mathbf{x}_{0}\leqslant\mathbf{x}_{1}$, set $\mathbf{y}_{0}:=A\mathbf{x}_{0}$ and $\mathbf{y}_{1}:=A\mathbf{x}_{1}$ and assume that $\mathbf{y}_{0}\leqslant\mathbf{y}_{1}$. Given $\mathbf{y}$ in the range of $A$ with $\mathbf{y}_{0}\leqslant\mathbf{y}\leqslant\mathbf{y}_{1}$, is there a vector $\mathbf{x}\in\mathbb{R}^{n}$ with $A\mathbf{x}=\mathbf{y}$, and $\mathbf{x}_{0}\leqslant\mathbf{x}\leqslant\mathbf{x}_{1}$?
The answer is no. Let
@@
Observe that the matrix $A$ is invertible and hence surjective. Clearly, $A(0,0)^{\mathsf{T}}=(0,0)^{\mathsf{T}}$ and $A(3,1)^{\mathsf{T}}=(15,4)^{\mathsf{T}}$. Now, consider the vector $(10,3)\in\mathbb{R}^{2}$: clearly $(0,0)\leqslant(10,3)\leqslant(15,4)$ but the only solution of the equation $A\mathbf{x}=(10,3)^{\mathsf{T}}$ is $\mathbf{x}=(1,2)$ which does not lie between $(0,0)$ and $(3,1)$.
@@
For $\mathbf{x}_{0}=(0,0,0)^{\mathsf{T}}$ and $\mathbf{x}_{1}=(1,6,3)^{\mathsf{T}}$, set $\mathbf{y}_{0}:=A\mathbf{x}_{0}=(0,0)^{\mathsf{T}}$ and $\mathbf{y}_{1}:=A\mathbf{x}_{1}=(4,9)^{\mathsf{T}}$. Consider the vector $\mathbf{y}=(3,1)^{\mathsf{T}}$. Indeed,
@@
$0\leqslant 3-t\leqslant 1,\quad 0\leqslant 1-t\leqslant 6,\quad\text{and}\quad 0\leqslant t\leqslant 3$
Hence there is no vector $(x,y,z)\in\mathbb{R}^{3}$ such that $(0,0,0)\leqslant(x,y,z)\leqslant(1,6,3)$ and $(0,0)^{\mathsf{T}}=A(0,0,0)^{\mathsf{T}}\leqslant(3,1)^{\mathsf{T}}=A(x,y,z)^{\mathsf{T}}\leqslant(4,9)=A(1,6,3)^{\mathsf{T}}$.
This example gives a clue to a better question. Indeed, notice that if we start with the inequality
@@
So, let us rephrase our question once again, by focusing on the vectors in the range of $A$, instead of starting with vectors in the domain.
@@
Let $A$ be an $m\times n$ real matrix and let $\mathbf{y}_{0}$ and $\mathbf{y}_{1}\in\mathbb{R}^{m}$, both in the range of $A$ such that $\mathbf{y}_{0}\leqslant\mathbf{y}_{1}$. Given $\mathbf{y}$ in the range of $A$ with $\mathbf{y}_{0}\leqslant\mathbf{y}\leqslant\mathbf{y}_{1}$, is it true that there exist $\mathbf{x}_{0}$, $\mathbf{x}$ and $\mathbf{x}_{1}\in\mathbb{R}^{n}$ with $A\mathbf{x}_{0}=\mathbf{y}_{0}$, $A\mathbf{x}=\mathbf{y}$, and $A\mathbf{x}_{1}=\mathbf{y}_{1}$, and such that $\mathbf{x}_{0}\leqslant\mathbf{x}\leqslant\mathbf{x}_{1}$?
In the next section, we attempt to answer Question 4.
3. Linear Transformations: Answers.
Before trying to answer Question 4, we need to make a couple of observations. The reader will have noticed that $\mathbf{x}_{0}=\mathbf{0}$ in all the examples in the previous section. The reason for that should be clear after the following two propositions. In the first one, we show that when considering Question 4 above for only one inequality, we may assume one side of the inequality equals zero. In the second proposition, we will show that this is in fact equivalent to the inequalities in Question 4.
@@
To prove that (1) implies (2), assume $\mathbf{0}\leqslant A\mathbf{x}$. Then, $A\mathbf{0}\leqslant A\mathbf{x}$. But (1) implies that there exist vectors $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$ such that $A\mathbf{z}_{1}=A\mathbf{0}=\mathbf{0}$, $A\mathbf{z}_{2}=A\mathbf{x}$ and $\mathbf{z}_{1}\leqslant\mathbf{z}_{2}$. Setting $\mathbf{x}^{\prime}:=\mathbf{z}_{2}-\mathbf{z}_{1}\geqslant\mathbf{0}$ we obtain the desired conclusion, since $A\mathbf{x}^{\prime}=A(\mathbf{z}_{2}-\mathbf{z}_{1})=A\mathbf{z}_{2}=A\mathbf{x}$ and $\mathbf{0}\leqslant\mathbf{x}^{\prime}$.
We now show that (2) implies (1). Assume $A\mathbf{x}_{1}\leqslant A\mathbf{x}_{2}$. Then we have that $A(\mathbf{x}_{2}-\mathbf{x}_{1})\geqslant\mathbf{0}$. But then (2) implies that there exists a vector $\mathbf{z}$ such that $A\mathbf{z}=A(\mathbf{x}_{2}-\mathbf{x}_{1})$ and $\mathbf{0}\leqslant\mathbf{z}$. Then, taking $\mathbf{x}_{2}^{\prime}:=\mathbf{x}_{1}+\mathbf{z}$ and $\mathbf{x}_{1}^{\prime}:=\mathbf{x}_{1}$ gives the result, since $\mathbf{x}_{2}^{\prime}-\mathbf{x}_{1}^{\prime}=\mathbf{z}\geqslant\mathbf{0}$, $A\mathbf{x}_{1}^{\prime}=A\mathbf{x}_{1}$, and $A\mathbf{x}_{2}^{\prime}=A(\mathbf{x}_{1}+\mathbf{z})=A\mathbf{x}_{2}$, as desired. ∎
It is natural to ask, and perhaps not obvious, if the conditions in the above proposition hold not only for two inequalities, but also for three. It turns out that they do hold.
@@
It seems that this definition was first given by Adi Ben-Israel and Thomas N.E. Greville in [2]. The full characterization of weakly monotone matrices is not easy and it is beyond the scope of this article (much more information can be found in [1, 10]). Nevertheless, we will give a partial characterization.
We start with a previous example. Let
@@
To verify that $A$ is weakly monotone, choose $\mathbf{x}=(x_{1},x_{2},x_{3})^{\mathsf{T}}$ such that $A\mathbf{x}\geqslant\mathbf{0}$. Then
@@
But clearly we can choose $\mathbf{x}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},x_{3}^{\prime})$ with $\mathbf{x}^{\prime}\geqslant\mathbf{0}$ with $x_{1}^{\prime}+x_{3}^{\prime}=x_{1}+x_{3}$ and with $x_{2}^{\prime}+x_{3}^{\prime}=x_{2}+x_{3}$ (for example, take $x_{1}^{\prime}:=x_{1}+x_{3}$, $x_{2}^{\prime}:=x_{2}+x_{3}$ and $x_{3}^{\prime}:=0$) and hence with $A\mathbf{x}=A\mathbf{x}^{\prime}$. Thus $A$ is indeed weakly monotone.
Before we continue talking about weakly monotone matrices, let us state the following definition which was originally given by Lothar Collatz [3] for square matrices.
@@
Collatz showed that a square matrix is monotone if and only if it is invertible and its inverse is nonnegative. [^4] The anxious reader can try and prove this now: it is not terribly difficult. The patient reader can prove this after seeing the results ahead.
Olvi L. Mangasarian [8] extended the above definition to nonsquare matrices and, furthermore, he showed that a matrix is monotone if and only if $A$ has a nonnegative left inverse. Of course, if $A$ has a left inverse, then $A$ is injective and, in particular, $m\geqslant n$.
In what follows, we intend to give a partial characterization of weakly monotone matrices, using the techniques in [10, 8]. We start with a famous lemma, whose proof we include here for the sake of completeness. The proof uses the hyperplane separating theorem for convex subsets of Euclidean space (see, for example, [7, p. 510]).
@@
does not contain the vector $\mathbf{b}$. Since $\mathcal{C}$ is a closed convex set, there exists $\mathbf{y}\in\mathbb{R}^{m}$ and $\alpha\in\mathbb{R}$ such that the hyperplane $\{\mathbf{w}\in\mathbb{R}^{m}\mid\mathbf{y}^{\mathsf{T}}\mathbf{w}=\alpha\}$ separates $\mathbf{b}$ from $\mathcal{C}$. In fact, $\mathbf{y}^{\mathsf{T}}b<\alpha$ and $\mathbf{y}^{T}\mathbf{w}>\alpha$ for all $\mathbf{w}\in\mathcal{C}$.
Observe that since $\mathbf{0}\in\mathcal{C}$, we have $\alpha<0$ and hence $\mathbf{y}^{\mathsf{T}}b<0$. For each $\lambda>0$ and each $k$, the vector $\lambda\mathbf{v}_{k}\in\mathcal{C}$. Therefore $\mathbf{y}^{\mathsf{T}}\lambda\mathbf{v}_{k}>\alpha$; equivalently, $\mathbf{y}^{\mathsf{T}}\mathbf{v}_{k}>\frac{\alpha}{\lambda}$. Since $\lambda>0$ can be arbitrarily large (and $\alpha<0$), this means that $\mathbf{y}^{\mathsf{T}}\mathbf{v}_{k}\geqslant 0$ for all $k$. But this is equivalent to $\mathbf{y}^{T}M\geqslant\mathbf{0}$. That is, (2) holds.
@@
Let $A$ be an $m\times n$ matrix of full rank. Assume that for all $\mathbf{x}\in\mathbb{R}^{n}$ with $A\mathbf{x}\geqslant\mathbf{0}$ there exists $\mathbf{x}^{\prime}\in\mathbb{R}^{m}$ such that $\mathbf{x}^{\prime}\geqslant\mathbf{0}$ and $A\mathbf{x}=A\mathbf{x}^{\prime}$. Then,
  • If $m\geqslant n$, there exists an $n\times m$ matrix $B$ with $BA=I_{n}$ such that $B\geqslant 0$.
  • If $m\leqslant n$, there exists an $n\times m$ matrix $B$ with $AB=I_{m}$ such that $B\geqslant 0$.